Run Meeting: August 22nd, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings

12:00PM RC Meeting (https://bluejeans.com/7576835804)

RC: S. Kuhn


2020

March
19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30,31
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21 22 23 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

Summary of Aug. 22th

Owl shift

  1. Shift summary https://logbooks.jlab.org/entry/3832186
  2. Issues: tunning the beam took a while and the beam was shifted. new target scan. new positions on 2c21, 2c24, and 2H01
  3. Recorded data: 0

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3832342
  2. Issues: None major (one run with 1 dead DREAM chip - fixed by recycling LV)
  3. Recorded data: 40 M triggers, incl. 12 M with no-roads trigger (rest standard trigger on D2)

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3832494
  2. Issues: DC HV (lost part of R2S6 during one run)
  3. Recorded data: about 4.4M events on D2 and about 30 M events on H2.

Whole day summary

  1. Total recorded data today: About 74 M.
  2. Part-II total accumulated data on D2: 773 M events
  3. Concerns: Continuing puzzle about the large backgrounds in BOM, downstream counters and all FD counters.


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded